Ragan Engines, Inc., was founded nine years ago by a brother and sister—Carrington and Genevieve Ragan—and has remained a privately owned company. The company manufactures marine engines for a variety of applications. Ragan has experienced rapid growth because of a proprietary technology that increases the fuel efficiency of its engines with very little sacrifice in performance. The company is equally owned by Carrington and Genevieve. The original agreement between the siblings gave each 125,000 shares of stock. Larissa has asked Dan to determine a value per share of Ragan stock. To accomplish this, Dan has gathered the following information about some of Ragan’s competitors that are publicly traded:

EPS DPS STOCK PRICE ROE 11.00% 14.00% 14.00 17.00 NA13.00 12.50% 14.67% Blue Ribband Motors Corp. $1.24 S.39 $20.10 Bon Voyage Marine, Inc. 155 47 Nautilus Marine Engines 6 3160 Industry average 16.85 S.85 $.51 $22.85

Nautilus Marine Engines’ negative earnings per share (EPS) was the result of an accounting write-off last year. Without the write-off, EPS for the company would have been $1.93. Last year, Ragan had an EPS of $3.65 and paid a dividend to Carrington and Genevieve of $195,000 each. The company also had a return on equity of 18 percent. Larissa tells Dan that a required return for Ragan of 13 percent is appropriate.
Q1. Assuming the company continues its current growth rate, what is the value per share of the company’s stock?
Q2. Dan has examined the company’s financial statements, as well as examining those of its competitors. Although Ragan currently has a technological advantage, Dan’s research indicates that Ragan’s competitors are investigating other methods to improve efficiency. Given this, Dan believes that Ragan’s technological advantage will last only for the next five years. After that period, the company’s growth will likely slow to the industry average. Additionally, Dan believes that the required return the company uses is too high. He believes the industry average required return is more appropriate. Under Dan’s assumptions, what is the estimated stock price?
Q3. What is the industry average price?earnings ratio? What is Ragan’s price?earnings ratio? Comment on any differences and explain why they may exist.

Q4. Assume the company’s growth rate slows to the industry average in five years. What future return on equity does this imply?
Q5. Carrington and Genevieve are not sure if they should sell the company. If they do not sell the company outright to East Coast Yachts, they would like to try and increase the value of the company’s stock. In this case, they want to retain control of the company and do not want to sell stock to outside investors. They also feel that the company’s debt is at a manageable level and do not want to borrow more money. What steps can they take to try and increase the price of the stock? Are there any conditions under which this strategy would not increase the stock price?
Answer #1

(1) Last EPS of Ragan Inc = $ 3.65 per share, Dividends Received = $ 195000 each to both partners, Stocks Held = 125000 each by both partners, ROE = 18 % and Required Rate of Return = r = 13 %

Dividend Per Share = D0 = 195000 / 125000 = $ 1.56 per share

Retention Ratio = [1-(1.56/3.65)] = 0.5726

ROE = 18 %

Sustainable Growth Rate = ROE x Retention Ratio = g = 0.18 x 0.5726 = 0.1031 or 10.31 %

Expected Dividend = D1 = D0 x (1+g) = 1.56 x 1.1031 = $ 1.721 approximately

Assuming that the company sutsains its current growth rate to perpetuity the current stock price P0 would be determined by the constant growth Gordon Growth Model:

P0 = D1 / (r-g) = 1.721 / (0.13-0.1031) = $ 63.98 approximately.

(2) Industry Average Figures:

DPS = $ 0.51 and EPS = $ 0.85

Retention Ratio =[1-(0.51 / 0.85)] = 0.4 and ROE = 12.5 %

Industry Average Growth = ROE x Retention Ratio = 0.125 x 0.4 = 0.05 or 5 %

Industry Average Required Rate of Return = R = 14.67 %

Ragan Engines Inc:

DPS0 = $ 1.56 and growth rate = 10.31 %

DPS1 = $ 1.721, DPS 2 = $ 1.898, DPS3 = $ 2.094, DPS4 = $ 2.3098, DPS5 = $ 2.548 and DPS6 = $ 2.811

The firm's DPS grows at the elevated rate of 10.31 % for five years owing to the company's technoological advantage, but slows down to the industry average of 5 % after five years. The discounting rate (required rate) to be used is the industry average rate of 14.67 %

Terminal Value of the perpetually growing dividend = 2.811 / (0.1467 - 0.05) = $ 29.069

PV of Terminal Value of perpetually growing dividend = 29.069 / (1.1467)^(5) = P1 = $ 14.66

PV of the elevated growth phase dividends = P2 = 1.721 / 1.1467 + 1.898 / (1.1467)^(2) + 2.094 / (1.1467)^(3) + 2.3098 / (1.1467)^(4) + 2.548 / (1.1467)^(5) = $ 6.955

Current Stock Price = P1 + P2 = 14.66 + 6.955 = $ 21.61

(3) Industry Average Figures:

DPS = $ 0.51 and EPS = $ 0.85

Retention Ratio =[1-(0.51 / 0.85)] = 0.4 and ROE = 12.5 %

Industry Average Growth = ROE x Retention Ratio = 0.125 x 0.4 = 0.05 or 5 %

Industry Average Required Rate of Return = R = 14.67 %

Expected Industry Average Dividend = D1 = 0.51 x 1.05 = $ 0.5355

Current Stock Price = P0 = D1 / (R-g) = 0.5355 / (0.1467 - 0.05) = $ 5.54

EPS = $ 0.85

Industry Average PE Ratio = P0 /EPS = 5.54 / 0.85 = 6.52

Ragan Stock Price = $ 63.98

EPS0 = $ 3.65

Ragan's Price PE Ratio = 63.98 / 3.65 = 17.53

The difference between the two PEs is owing to the difference between the type of stocks they are. Ragan Engine Inc has retention ratio higher than the industry average, which implies an above industry average growth rate and therefore classification as a growth stock. Growth stocks possess high PE ratios as their growing price is primarily attributable to their higher than average earnings capacity and their ability to successfully reinvest these earnings to into the business to generate even more growth. Hence, Ragan Inc's PE is higher than the industry average PE.

(4) Ragan Inc's Growth Rate = 10.31 %, Current EPS = $ 3.65, Current DPS = $ 1.56

EPS5 = 3.65 x (1.1031)^(5) = $ 5.962 and DPS5 = 1.56 x (1.1031)^(5) = $ 2.548

At end of Year 5, the growth slows down to the industry average of 5 % per annum and assuming that the retention ratio remains the same as before, 0.05 (Industry Average Growth) = New ROE x [1-(2.548/5.962)]

New ROE = 0.05 / 0.5726 = 0.0873 or 8.73 % approximately

NOTE: Please raise a separate qeury for solution to the last sub-part.